Go to the U of M home page
School of Physics & Astronomy
Physics & Astronomy Wiki

groups:homestake:meetings:20150305

Homestake meeting minutes, 03/05/15

Attending: Victor, Tanner, Noah, Vuk, Daniel, Pat

Homestake wiki: https://zzz.physics.umn.edu/groups/homestake/home

Our agenda:

1) Homestake status (Tanner, Gary, Daniel)

  • stations duty cycle: Tanner hasn't seen any interruptions in the data flow. A few weeks ago there was a strange issue: mass positions for 4100D for two days got lumped into one day file. Will have to deal with this scenario in the code that makes frames, working on implementing this. Added data quality from the surface Q330, so that we can check the clock quality there for diagnostic purposes later.
  • data flow: Daniel: data coming in, trying to convert antelope files into a different format for analysis. Tanner: data is being forwarded to IRIS, started last week. Victor: there may be something wrong with the instrument response information (spectra appear to be 1000x too high), will have to follow up on this.
  • frames: Tanner making frames (mid February now, trying to catch up)
  • data quality shifts: Tanner working on the data quality monitor, need some time for it.

2) Planning next visit (Mar 23 - Apr 3)

  • 2000 level: schematic Vuk had a discussion with Bryce et al. They are currently installing the fiber at the 2000 level. Currently planning 3 stations there, all northwest from the existing stations (i.e. northwest from the yellow triangle in the linked schematic). Two of the stations will require batteries, the third already has a concrete pad, power and network. They are already installing the network cable, should be done this week.
  • locations, foam panels, granite tiles, network, GPS: Vuk asked for 15 foam panels to be moved to the 2000 level. Should also have enough granite tiles, but Vuk found a local vendor in Rapid City that carries them. Jamey installed another splitter in the GPS distribution system that will cover the 2000 level, Tanner confirmed it works well.
  • batteries, chargers: Terry putting together the diode boxes. The power at the 2000 level will not be sufficient to recharge 20 batteries overnight. Looking at another solution where the batteries could be charged slower (2-3 days). Bryce offered to help with these, they could visit the 2000 level once a month and swap the batteries for us.
  • radios: purchased mast, should be delivered soon.
  • Victor: in July/August we could have a student from Lee Liberty for active seismic measurements. Victor will follow up in April.
  • Vuk: we will also have to borrow Octans. Will contact PASSCAL about this after the next trip.

3) Analysis of existing data

  • radiometer approach (Noah, Tanner): pdf
groups/homestake/meetings/20150305.txt · Last modified: 2015/03/05 16:42 by mandic